In a reaction, 21.2 g of chromium(III) oxide reacts with 7.53 g of aluminum to produce chromium and aluminum oxide. If 14.5 g of chromium is produced, what mass of aluminum oxide is produced?
Answer #1

Solution-

THe reaction can be written as

Cr2O3 + 2Al ----> 2Cr + Al2O3

The molar mass of Cr2O3 is = (2*52) +(3*16)
= 152 g /mol
The molar mass of Al is
= 27 g / mol
The molar mass of Al2O3 is = (2*27)+(3*16)
= 102 g/mol
The molar mass of Cr is = 52 g/mol

Now from the balanced Equation

152 g of Cr2O3 reacts = 2*27=54 g of Al

So,
21.2 g of Cr2O3 reacts with Y g of Al
Y = ( 21.2*54) / 152
= 7.53 g of Al

Hence both the two reagents are completely reacted

152 g of Cr2O3 produces 102 g of Al2O3
30.2 g of Cr2O3 produces X g of Al2O3
X = ( 21.2*102) / 152
= 14.23 g of Al2O3